# Out-of-Distribution Detection in Molecular Complexes via Diffusion Models for Irregular Graphs

### v1

@@ 5.9 Acknowledgments
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Relation to score error.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
S1.7.2 PF-ODE negative log-likelihood as a random variable
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
S1.7.3 Coupling PF-ODE likelihoods to GEMS errors
@@
Assumption S1.8 states that for any two complexes $x_{1},x_{2}$ with $L(x_{1})\leq L(x_{2})$ (i.e. $x_{1}$ is at least as likely as $x_{2}$ under the diffusion model), the worst-case prediction error at $x_{1}$ is no larger than the worst-case error at $x_{2}$. The function $\phi$ acts as a calibration curve mapping PF-ODE NLL values to an upper bound on the predictive error. In practice, $\phi$ can be obtained empirically by fitting a non-decreasing upper envelope (e.g. a scaled and shifted exponential) to the scatter plot of $(L(x),e_{\theta}(x))$ on a held-out calibration set (Figure S1). This is in accordance with the further empirical findings explained in Section 2.4; for visualizations, refer to Figures 3 and 4 in the main text.
@@
[…]
@@
[…]
@@
[…]
@@
S1.8 Comprehensive Error Tables
@@
Table S2: Embedding-space GEMS baseline OOD detection performance across held-out splits. For each dataset, we report the mean $\pm$ standard deviation of AUROC, accuracy (Acc), F1, and specificity (Spec) over balanced bootstrap test subsets, treating OOD as the positive class. For each split we take the best detector by AUROC among Mahalanobis, $k$-NN distance, Local Outlier Factor, Isolation Forest and one-class SVM. The final rows report the mean and standard deviation across datasets for these summary metrics. Per-dataset precision (Prec) and recall of the embedding-space GEMS baseline, using the best detector per dataset selected by AUROC. We report mean $\pm$ standard deviation over balanced bootstrap test subsets, with OOD as the positive class.
@@
Table S3: Performance of the trajectory-aware LDR OOD detector across held-out splits: For each OOD dataset, we report the mean $\pm$ standard deviation of AUROC, accuracy (Acc), F1, specificity (Spec), Precision (Prec) and recall over balanced bootstrap test subsets, treating OOD as the positive class.
The embedding-based baseline achieved a respectable mean AUROC of 0.78 and F1-score of 0.73 across all datasets. However, performance was highly inconsistent. While it effectively detected anomalies in some distinct OOD splits (e.g., HIV proteases), it degraded to near-random guessing on challenging datasets with moderate distribution shift like CASF2016. This baseline serves as a strong and interpretable reference for our experiments, demonstrating what is achievable using post-hoc analysis of fixed representations. Overall, it systematically underperformed the trajectory–aware OOD classifier, which leverages multiple trajectory statistics derived from our diffusion model.
Dataset AUROC Acc F1 Spec Prec Recall
[…]
Table S4: Rate-In OOD Detection Performance: For each OOD dataset, we report the mean $\pm$ standard deviation of AUROC, accuracy (Acc), F1, and specificity (Spec) over 100 bootstrap iterations with 70/30 train/test splits, treating OOD as the positive class. The final rows report the mean and standard deviation across datasets for these summary metrics. For each OOD dataset, we report precision and recall over 100 bootstrap iterations.

### v2

@@ 4.9 Acknowledgments
$\frac{\mathrm{d}}{\mathrm{d}t}D_{t}$ $=\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}p_{t}(x)\,\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x$ (S1.8)
$=\int_{\mathbb{R}^{d}}(\partial_{t}p_{t}(x))\,\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x\;+\;\int_{\mathbb{R}^{d}}p_{t}(x)\,\partial_{t}\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x.$ (S1.9)
@@
$\int p_{t}\,\partial_{t}\log\!\frac{p_{t}}{q_{t}}\,\mathrm{d}x$ $=\int\partial_{t}p_{t}(x)\,\mathrm{d}x\;-\;\int\frac{p_{t}(x)}{q_{t}(x)}\,\partial_{t}q_{t}(x)\,\mathrm{d}x.$ (S1.11)
Because $p_{t}$ is a probability density, $\int\partial_{t}p_{t}\,\mathrm{d}x=\partial_{t}\int p_{t}\,\mathrm{d}x=0$, hence
$\frac{\mathrm{d}}{\mathrm{d}t}D_{t}=\int(\partial_{t}p_{t})\,\log\!\frac{p_{t}}{q_{t}}\,\mathrm{d}x\;-\;\int\frac{p_{t}}{q_{t}}\,\partial_{t}q_{t}\,\mathrm{d}x.$ (S1.12)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}D_{t}$ $=-\int\nabla_{x}\cdot(p_{t}v^{\star})\,\log\!\frac{p_{t}}{q_{t}}\,\mathrm{d}x\;+\;\int\frac{p_{t}}{q_{t}}\,\nabla_{x}\cdot(q_{t}v_{\phi})\,\mathrm{d}x.$ (S1.14)
@@
$-\int\nabla_{x}\cdot(p_{t}v^{\star})\,\log\!\frac{p_{t}}{q_{t}}\,\mathrm{d}x$ $=\int p_{t}(x)\,v^{\star}(x,t)\cdot\nabla_{x}\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x.$ (S1.15)
@@
$\int\frac{p_{t}}{q_{t}}\,\nabla_{x}\cdot(q_{t}v_{\phi})\,\mathrm{d}x$ $=-\int q_{t}(x)\,v_{\phi}(x,t)\cdot\nabla_{x}\!\left(\frac{p_{t}(x)}{q_{t}(x)}\right)\,\mathrm{d}x$ (S1.16)
$=-\int p_{t}(x)\,v_{\phi}(x,t)\cdot\nabla_{x}\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x.$ (S1.17)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}D_{t}=\int p_{t}(x)\,\bigl(v^{\star}(x,t)-v_{\phi}(x,t)\bigr)\cdot\nabla_{x}\log\!\frac{p_{t}(x)}{q_{t}(x)}\,\mathrm{d}x,$ (S1.18)
@@
$\bigl|\dot{D}_{t}\bigr|\;\leq\;\left(\int_{\mathbb{R}^{d}}p_{t}(x)\,\bigl\|v^{\star}(x,t)-v_{\phi}(x,t)\bigr\|_{2}^{2}\,\mathrm{d}x\right)^{\!1/2}\sqrt{I_{t}},$ (S1.19)
@@
Relation to score error
@@
$A_{t}=\frac{g(t)^{4}}{4}\int_{\mathbb{R}^{d}}p_{t}(x)\,\bigl\|s^{\star}(x,t)-s_{\phi}(x,t)\bigr\|_{2}^{2}\,\mathrm{d}x.$ (S1.23)
@@
$\mathcal{L}_{\mathrm{score}}(\phi)\;:=\;\int_{0}^{1}g(t)^{4}\int_{\mathbb{R}^{d}}p_{t}(x)\,\bigl\|s^{\star}(x,t)-s_{\phi}(x,t)\bigr\|_{2}^{2}\,\mathrm{d}x\,\mathrm{d}t,$ (S1.24)
@@
$D_{0}\;=\;-\int_{0}^{1}\dot{D}_{t}\,\mathrm{d}t\;\leq\;\int_{0}^{1}\bigl|\dot{D}_{t}\bigr|\,\mathrm{d}t\;\leq\;\int_{0}^{1}\sqrt{A_{t}I_{t}}\,\mathrm{d}t.$ (S1.26)
@@
$D_{0}\;\leq\;\left(\int_{0}^{1}A_{t}\,\mathrm{d}t\right)^{1/2}\left(\int_{0}^{1}I_{t}\,\mathrm{d}t\right)^{1/2}\;=\;\frac{1}{2}\,\sqrt{\mathcal{L}_{\mathrm{score}}(\phi)}\left(\int_{0}^{1}I_{t}\,\mathrm{d}t\right)^{1/2}.$ (S1.27)
@@
S1.6.2 PF-ODE negative log-likelihood as a random variable
@@
$\mathrm{KL}(p_{0}\|q_{0})$ $=\int_{\mathbb{R}^{d}}p_{0}(x)\,\log\!\frac{p_{0}(x)}{q_{0}(x)}\,\mathrm{d}x$ (S1.32)
$=\int_{\mathbb{R}^{d}}p_{0}(x)\,\log p_{0}(x)\,\mathrm{d}x-\int_{\mathbb{R}^{d}}p_{0}(x)\,\log q_{0}(x)\,\mathrm{d}x$ (S1.33)
$=-H(p_{0})+\int_{\mathbb{R}^{d}}p_{0}(x)\,L(x)\,\mathrm{d}x.$ (S1.34)
@@
$\mathrm{Var}_{x\sim p_{0}}\bigl(L(x)\bigr)\;=\;\int_{\mathbb{R}^{d}}p_{0}(x)\,\bigl(L(x)-L_{\mathrm{typ}}\bigr)^{2}\,\mathrm{d}x\;\leq\;\sigma^{2}.$ (S1.37)
@@
$\int_{\{x:\,L(x)\,\geq\,L_{\mathrm{typ}}+\alpha\}}p_{0}(x)\,\mathrm{d}x\;\leq\;\frac{\sigma^{2}}{\alpha^{2}}.$ (S1.38)
@@
$\int_{\{x:\,|Z(x)|\,\geq\,\alpha\}}p_{0}(x)\,\mathrm{d}x\;\leq\;\frac{\mathrm{Var}_{p_{0}}(Z)}{\alpha^{2}}\;\leq\;\frac{\sigma^{2}}{\alpha^{2}}.$ (S1.39)
@@
S1.6.3 Coupling PF-ODE likelihoods to GEMS errors
@@
Assumption S1.8 states that for any two complexes $x_{1},x_{2}$ with $L(x_{1})\leq L(x_{2})$ (i.e. $x_{1}$ is at least as likely as $x_{2}$ under the diffusion model), the worst-case prediction error at $x_{1}$ is no larger than the worst-case error at $x_{2}$. The function $\phi$ acts as a calibration curve mapping PF-ODE NLL values to an upper bound on the predictive error. In practice, $\phi$ can be obtained empirically by fitting a non-decreasing upper envelope (e.g. a scaled and shifted exponential) to the scatter plot of $(L(x),e_{\theta}(x))$ on a held-out calibration set (Figure S1). This is in accordance with the further empirical findings explained in Section 2.4; for visualizations, refer to Figures 5 and 6 in the main text.
@@
$\int_{\{x:\,e_{\theta}(x)\,>\,\phi(L_{\mathrm{typ}}+\alpha)\}}p_{0}(x)\,\mathrm{d}x\;\leq\;\frac{\sigma^{2}}{\alpha^{2}}.$ (S1.41)
@@
$\int_{\{x:\,e_{\theta}(x)\,>\,\phi(L_{\mathrm{typ}}+\alpha)\}}p_{0}(x)\,\mathrm{d}x\;\leq\;\int_{A_{\alpha}^{c}}p_{0}(x)\,\mathrm{d}x=\int_{\{x:\,L(x)\,>\,L_{\mathrm{typ}}+\alpha\}}p_{0}(x)\,\mathrm{d}x.$ (S1.45)
@@
$\int_{\{x:\,L(x)\,>\,L_{\mathrm{typ}}+\alpha\}}p_{0}(x)\,\mathrm{d}x\;\leq\;\frac{\sigma^{2}}{\alpha^{2}},$ (S1.46)
@@
S1.7 Comprehensive Error Tables
To facilitate a rigorous comparison across our diffusion trajectory-based OOD detection and the implemented baselines Rate-In and Embedding Space, we present detailed performance metrics for each approach. These comprehensive error tables provide an overview over the performance on the OOD datasets employed in this work, reporting the mean and standard deviation for AUROC, accuracy, F1-score, specificity, precision, and recall. By evaluating our diffusion trajectory approach alongside the Embedding space and Rate-In baselines, we can assess not only the raw discriminative power (AUROC) but also the practical reliability and classification balance of each method. All results are derived using the balanced bootstrap evaluation strategy to ensure that the reported statistics are resilient to the inherent class imbalances present in the protein-family-based OOD splits.
@@
Table S1: Embedding space GEMS baseline OOD detection performance across held-out splits: For each dataset, we report the mean $\pm$ standard deviation of AUROC, accuracy (Acc), F1, and specificity (Spec) over balanced bootstrap test subsets, treating OOD as the positive class. For each split we take the best detector by AUROC among Mahalanobis, $k$-NN distance, Local Outlier Factor, Isolation Forest and one-class SVM. The final rows report the mean and standard deviation across datasets for these summary metrics. Per-dataset precision (Prec) and recall of the embedding-space GEMS baseline, using the best detector per dataset selected by AUROC. We report mean $\pm$ standard deviation over balanced bootstrap test subsets, with OOD as the positive class.
Dataset AUROC Acc F1 Spec Prec Recall
[…]
3dd0 $0.337\,{\pm}\,0.029$ $0.498\,{\pm}\,0.005$ $0.238\,{\pm}\,0.317$ $0.641\,{\pm}\,0.478$ $0.179\,{\pm}\,0.239$ $0.354\,{\pm}\,0.472$
Table S2: Rate-In OOD detection performance across held-out splits: For each OOD dataset, we report the mean $\pm$ standard deviation of AUROC, accuracy (Acc), F1, and specificity (Spec) over 100 bootstrap iterations with 70/30 train/test splits, treating OOD as the positive class. The final rows report the mean and standard deviation across datasets for these summary metrics. For each OOD dataset, we report precision and recall over 100 bootstrap iterations.
@@
Table S3: Performance of the diffusion trajectory LDR OOD detector across held-out splits: For each OOD dataset, we report the mean $\pm$ standard deviation of AUROC, accuracy (Acc), F1, specificity (Spec), Precision (Prec) and recall over balanced bootstrap test subsets, treating OOD as the positive class.